OpenStudy (issy14):

Derive: (I'm lost on this one, I'm not sure which one is f(x) and which one is g(x). We are suppose to apply the chain rule here: f'(g(x)) g(x). w= 100e^-x^2

OpenStudy (precal):

I would just take the derivative of both sides so it becomes w ' =

OpenStudy (precal):

Let me type the equation first to verify

OpenStudy (precal):

\[w=100e ^{-x ^{2}}\]

OpenStudy (precal):

is this correct before we do the chain rule

OpenStudy (issy14):

yes correct

OpenStudy (precal):

the derivative of e^u is u ' (e^u)

OpenStudy (precal):

wait your notation is throwing me off, let me apply chain rule and then explain it

OpenStudy (issy14):

ok

OpenStudy (precal):

u=-x^2 u ' = -2x

OpenStudy (precal):

\[w ' =-2x(100)e ^{-x^2}\]

OpenStudy (precal):

\[w ' =-200e ^{-x^2}\]

OpenStudy (precal):

ok that is your answer you need to memorize d/dx of e^u= u ' (e^u) this way your chain rule is built in

OpenStudy (precal):

hope that helps you

OpenStudy (issy14):

ok, may I ask you another question, Is that how I should approach chain rule problems?

OpenStudy (precal):

inner function by the way is u so that is your g(x) and the outer function is f(x) in this case 100e^x I find it easier to say who is the inner function, call it u and then take the derivative of u and call it u '

OpenStudy (issy14):

I always try to separate it as well, this one just threw me off completely thank you.

OpenStudy (issy14):

I'm taking calc I over the summer and its not an easy task